# Efficient Rule Induction by Ignoring Pointless Rules

### v2

@@ 5 Experiments
To answer Q4, we compare Reducer against Popper, Aleph (Srinivasan 2001), and ASPSynth (Bembenek, Greenberg, and Chong 2023).
@@
IGGP. In inductive general game playing (IGGP) (Cropper, Evans, and Law 2020), the task is to induce rules from game traces from the general game playing competition (Genesereth and Björnsson 2013).
IMDB. We use a real-world dataset which contains relations between movies, actors, and directors (Mihalkova, Huynh, and Mooney 2007).
@@ Results
Figure 2 shows the difference in predictive accuracy of Reducer vs Popper. Significance tests confirm $(p<0.05)$ that Reducer increases accuracy on 6/449 (1%) tasks. It decreases accuracy on 5/449 (1%) tasks. There is no significant difference in the other tasks.
@@ 6 Conclusions and Limitations
We have introduced an approach that identifies and ignores pointless (reducible and indiscriminate) rules. We have shown that ignoring pointless rules is optimally sound (Propositions 2 and 4). We implemented our approach in Reducer, which identifies pointless rules in hypotheses and builds constraints from them to prune the hypothesis space. We have proven that Reducer always learns an optimal hypothesis if one exists (Theorem 1). We have experimentally shown on multiple domains, including visual reasoning and game playing, that our approach can reduce learning times by 99% whilst maintaining high predictive accuracies. Furthermore, a key contribution of this paper is demonstrating that a conceptually straightforward yet previously unexplored idea (identifying pointless rules in hypotheses) can lead to substantial performance gains.
@@
The benefits of ignoring pointless rules should generalise to other ILP approaches. For instance, the improvements should directly improve hopper (Purgał, Cerna, and Kaliszyk 2022), which learns higher-order programs, and propper (Hillerström and Burghouts 2025), which uses neurosymbolic inference to learn programs from probabilistic data. Future work should empirically show how our idea benefits these systems.

### v3

@@ 5 Experiments
To answer Q4, we compare Reducer against Popper, Aleph (Srinivasan 2001), and ASPSynth (Bembenek et al. 2023).
@@
Reproducibility.
The code and experimental data for reproducing the experiments are available as supplementary material and will be made publicly available if the paper is accepted for publication.
@@
IGGP. In inductive general game playing (IGGP) (Cropper et al. 2020), the task is to induce rules from game traces from the general game playing competition (Genesereth and Björnsson 2013).
IMDB. We use a real-world dataset which contains relations between movies, actors, and directors (Mihalkova et al. 2007).
@@ Results
Figure 2 shows the difference in predictive accuracy of Reducer vs Popper. Significance tests confirm $(p<0.05)$ that Reducer increases accuracy on 6/449 (1%) tasks. It decreases accuracy on 5/449 (1%) tasks. There is no significant difference in the other tasks. The slight accuracy discrepancy is because multiple optimal hypotheses may exist. The two systems may find different ones. Two hypotheses with the same training cost might have different test accuracy.
@@ 6 Conclusions and Limitations
We have introduced an approach that identifies and ignores pointless (reducible and indiscriminate) rules. We have shown that ignoring pointless rules is optimally sound (Propositions 2 and 4). We implemented our approach in Reducer, which identifies pointless rules in hypotheses and builds constraints from them to prune the hypothesis space. We have proven that Reducer always learns an optimal hypothesis if one exists (Theorem 1). We have experimentally shown on multiple domains, including visual reasoning and game playing, that our approach can reduce learning times by 99% whilst maintaining high predictive accuracies.
@@
The benefits of ignoring pointless rules should generalise to other ILP approaches. For instance, the improvements should directly improve hopper (Purgał et al. 2022), which learns higher-order programs, and propper (Hillerström and Burghouts 2025), which uses neurosymbolic inference to learn programs from probabilistic data. Future work should empirically show how our idea benefits these systems.